Perpendicular to the graph of y = 1/3 x + 7; y-intercept (0,5)
Answer by nerdybill(7003)


Perpendicular to the graph of y = 1/3 x + 7; y-intercept (0,5)
.56
Because
y = 1/3 x + 7
is in "slope-intercept" form: y=mx+b
We can see that the slope is: 1/3
.
For a line to be perpendicular, the slope of the new line must be the "negative reciprocal":
(1/3)m = -1
m = -3 (slope of our new line)
.
Given that the problem stated that the y-intercept is at (0,5) this means that the 'b' (in the slope intercept form) is 5
.
So, our equation is:
y = -3x + 5 (this is what they're looking for)